### v2

# Security and Privacy on Generative Data in AIGC: A Survey

Abstract
The advent of artificial intelligence-generated content (AIGC) represents a pivotal moment in the evolution of information technology. With AIGC, it can be effortless to generate high-quality data that is challenging for the public to distinguish. Nevertheless, the proliferation of generative data across cyberspace brings security and privacy issues, including privacy leakages of individuals and media forgery for fraudulent purposes. Consequently, both academia and industry begin to emphasize the trustworthiness of generative data, successively providing a series of countermeasures for security and privacy. In this survey, we systematically review the security and privacy on generative data in AIGC, particularly for the first time analyzing them from the perspective of information security properties. Specifically, we reveal the successful experiences of state-of-the-art countermeasures in terms of the foundational properties of privacy, controllability, authenticity, and compliance, respectively. Finally, we summarize the open challenges and potential exploration directions from each of theses properties.
Index Terms:
Information security, AIGC, generative data, privacy, controllability, authenticity, compliance.
I Introduction
I-A Background
Artificial intelligence-generated content (AIGC) emerges as a novel generation paradigm for the production, manipulation, and modification of data. It utilizes advanced artificial intelligence (AI) technologies to automatically generate high-quality data at a rapid pace, including images, videos, text, audio, and graphics. With the powerful generative ability, AIGC can save time and unleash creativity, which are often challenging to achieve with professionally generated content (PGC) and user-generated content (UGC). Such progress in data creation can drive the emergence of innovative industries, particularly Metaverse [1], where digital and physical worlds converge. Early AIGC is limited by the algorithmic efficiency, hardware performance, and data scale, hindering the ability to fulfill optimal creation tasks. With the iterative updates of generative structures, notably generative adversarial networks (GANs) [2], AIGC has witnessed significant breakthroughs, generating realistic data that is often indistinguishable by humans from real data. In the generation of visual content, NVIDIA released StyleGAN [3] in 2018, which enables the controllable generation of high-resolution images and has undergone several upgrades. The subsequent year, DeepMind released DVD-GAN [4], which is designed for continuous video generation and exhibits great efficacy in complex data domains. Recently, diffusion models (DMs) [5] show more refined and novel image generation via the incremental noise addition. Guided by language models, DMs can improve the semantic coherence between input prompts and generated images. Excellent diffusion-based products, e.g., Stable Diffusion [^1] https://stability.ai/stablediffusion, Midjourney [^2] https://www.midjourney.com/, and Make-A-Video [^3] https://makeavideo.studio/, are capable of generating visually realistic images or videos that meet the requirements of diverse textual prompts. In the generation of language content, more attention is focused on ChatGPT, which reached 1.76 billion visits in May 2023. Trained on a large-scale text dataset, ChatGPT exhibits impressive performance in various contexts, including human-computer interaction and dialogues. For instance, researchers released LeanDojo [6], an open-source mathematical proof platform based on ChatGPT, providing toolkits, benchmarks, and models to tackle complex proof of formulas in an interactive environment. The integration of ChatGPT into the Bing search engine enhances search experiences, enabling users to effortlessly access comprehensive information. This powerful multi-purpose adaptability further exemplify the possibilities for humanity to achieve artificial general intelligence (AGI). Overall, compared to the PGC and UGC, AIGC demonstrates more advantages in data creation. AIGC possesses the ability to swiftly produce high-quality content while catering to personalized demands from users. As AI technology continues to advancements, the generative capability of AIGC is growing rapidly, promoting increased social productivity and economic value.
I-B Motivation
A large amount of generative data floods cyberspace, further enriching the diversity and abundance of online content. These generative data encompass multimodal information, which can be observed in various domains, e.g., news reporting, computer games, and social sharing. According to the Gartner’s report, AIGC will be anticipated to account for over 10% of all data creation in 2025 [7]. However, the proliferation of generative data also poses security and privacy issues. Firstly, generative data can expose individual privacy content by replicating training data. Generative models rely on large-scale data, which includes private content, e.g., faces, addresses, and emails. Existing works have demonstrated the memorization capabilities of large generative models [8, 9], leading to the potential replication of all or parts of the training data. This means that the generative data may also contain sensitive content which present in the training data. With specific prompts, GPT-2 can output personal information, including the name, the address, and the e-mail address [8]. An alarming study [9] revealed that Google’s Imagen can be prompted to output real-person photos, posing a significant threat to individual privacy. Therefore, it is necessary to hinder the generation of data containing privacy content. Secondly, generative data used for malicious purposes often involves false information, which can deceive the public, posing potential threats to both society and individuals. Recently, a false tweet about an explosion near the Pentagon went viral on social media, fooling many authoritative media sources and triggering fluctuations in the US stock market [10]. Moreover, the mature DeepFake technologies allow for the creation of convincing fake personal videos, which can be used to maliciously fabricate celebrity events [11]. The difficulty in discerning authenticity exposes the public to believing such content, resulting in severe damage to the reputations of celebrities. Thus, it is important to provide effective technology to confirm the authenticity of generative data. Meanwhile, generative data is required to have the controllability so that such potential threats can be proactively prevented. Thirdly, regulators around the world have further requirements for the compliance of generative data due to the critical implications of AIGC. Data protection regulatory authorities in Italy, Germany, France, Spain, and other countries have expressed concerns and initiated investigations into AIGC. In particular, China has taken a significant step by introducing the interim regulation on the management of generative artificial intelligence (AI) services [12]. This regulation encourages innovation in AIGC while mandating that generative data is non-toxic and factual. To adhere to the relevant regulations, it becomes crucial to ensure the compliance of generative data.
TABLE I: Comparisons of our work with existing surveys.
Reference Privacy Controllability Authenticity Compliance
Privacy in AIGC AIGC for Privacy Access Control Traceability Generative Detection Generative Attribution Non-toxicity Factuality
[13]
[14]
[15]
Ours
I-C Comparisons with Existing Surveys
Several works [13, 14, 15] have investigated the security and privacy in AIGC from different perspectives. Wang et al. [13] presents an in-depth survey of AIGC working principles, and roundly explored the taxonomy of security and privacy threats to AIGC. Meanwhile, they extensively reviewed solutions for intellectual property (IP) protection for AIGC models and generative data, with a focus on watermarking. Yet, they fail to provide countermeasures for other threats such as the utilization of non-compliant data.
Chen et al. [14] discussed three main concerns for promoting responsible AIGC, including 1) privacy, 2) bias, toxicity, misinformation, and 3) intellectual property. They summarized the issues and listed solutions related to existing AIGC products, e.g., ChatGPT, Midjourney, and Imagen. Nevertheless, they overlooked the importance of considering the authenticity of generative data in responsible AIGC.
Chen et al. [15] summarized the AIGC technology and analyzed the security and privacy challenges in AIGC. Moreover, they explored the potential countermeasures with advanced technologies, involving privacy computing, blockchain, and beyond. However, they did not pay attention to the detection and access control of generative data.
Figure 1: The process of AIGC. Real data collected is used to train generative models. Then generative models produce generative data. Finally, generative data are further analyzed. For generative data, there are corresponding protection requirements of security and privacy at different stages, which can be divided into privacy, controllability, authenticity, and compliance.
The differences between our work and previous works are: (1) Our work is targeted at generative data rather than AIGC. Previous works also explored issues about privacy in data collection and security of models. Yet, these issues are generic to AI security, which has been discussed in some works [16, 17]. (2) Previous works presented the corresponding techniques in terms of specific issues but neglected the fact that the same technique can address multiple issues. For example, watermarking can be used to protect copyright while detecting false data. On the contrary, we discuss security and privacy from the fundamental properties of information security, whose presence are ensured through specific techniques. (3) We supplement security issues that are not discussed in previous works, including access control and generative detection. In addition, we explore the use of generative data to power the privacy protection of real data. Table I shows the comparisons of our work with existing surveys.
In brief, the main contributions of our work are as follows:
  • We investigate the security and privacy issues on generative data in AIGC and comprehensively survey the corresponding state-of-the-art countermeasures.
  • We discuss security and privacy from a new perspective, i.e., the fundamental properties of information security, including privacy, controllability, authenticity, and compliance.
  • We point out the valuable future directions in security and privacy, toward building trustworthy generative data.
The rest of the paper is organized as follows. Section II reviews the basic AIGC process and categorizes the security and privacy on generative data in AIGC. In Sections III to VI, we discuss the issues and review the corresponding solutions from the perspectives of privacy, controllability, authenticity, and compliance, respectively. We present our outlook and suggest some future directions in Section VII. Finally, we summarize our work in Section VIII.
II Overview
II-A Process of AIGC
As illustrated in Fig. 1, we first discuss the AIGC process as follows:
II-A1 Real Data for Training
The data used for training impacts the features and patterns learned by AIGC models. Therefore, high-quality data forms the cornerstone of AIGC technology. Data collection typically involves various open-source repositories, including public databases, social media platforms, and online forums. These diverse sources provide AIGC training with a large-scale and diverse dataset. After collection, data filtering is applied to ensure the data quality, which involves removing irrelevant data and balancing the dataset for unbiased training. Additionally, data preprocessing, data augmentation, and data privacy protection steps can be undertaken based on different tasks to further enhance the quality and security of the training data.
II-A2 Generative Model in Training
The obtained data is used to train generative models, which are often performed by a centralized server with powerful computational capabilities. During training, generative models learn patterns and features in the data to generate results with a similar distribution to real data. Popular generative model architectures include generative adversarial networks (GANs), variational autoencoders (VAEs), flow-based models (Flows), and diffusion models (DMs), each with its strengths and weaknesses. The choice of models depends on specific requirements of tasks, available data, and computational resources. It is also important to note that training generative models requires substantial computational resources. On this basis, model fine-tuning is the process of adapting the pre-trained large model to a new task or domain without retraining. It only adjusts model parameters by training appropriate amount of additional data.
II-A3 Generative Data
After generative models are trained, they can be utilized to produce data. During this stage, users typically provide an input condition, e.g., a question or a piece of text. Then the model starts outputting data based on the input condition. In the generation of language content, AIGC exhibits the capability to outpace human authors in rapidly generating high-quality text, e.g., codes and articles. Additionally, it can engage in conversational interactions akin to humans, assisting users with various tasks and inquiries. The efficiency of AIGC in content creation and human-like interactions revolutionize how information is produced and communicated. In the generation of visual content, AIGC harnesses the powerful generative capabilities of models like DMs, enabling the generation of new images with realistic quality. Moreover, AIGC holds potential for video generation, as they can simultaneously process multiple video frames automatically.
II-A4 Analysis for Generative Data
After data generation, further analysis of the generative data is necessary to ensure the quality of generative data.
Generative data needs to undergo a quality assessment to check its accuracy, consistency, and integrality. If the generative data lacks in certain aspects, it requires model adjustments to improve the quality of generative data. Additionally, analyzing the risks associated with generative data can identify potential hazards. For instance, it is required to analyze whether there is discriminatory content, false information, or misleading content. By promptly detecting and addressing these issues, the negative impact of generative data can be minimized.
II-B Security and Privacy on Generative Data
Figure 2: The subclassification of security and privacy on generative data.
For generative data, there are corresponding security and privacy requirements at different stages. As shown in Fig. 1, we categorize these requirements according to the fundamental properties of information security, including privacy, controllability, authenticity, and compliance. Additionally, Fig. 2 shows the further subclassification.
II-B1 Privacy
Privacy refers to ensuring that individual sensitive information is protected. Generative data mimics the distribution of real data, which brings negative and positive impacts on the privacy of real data. Specifically, the following two impacts exist:
  • Privacy in AIGC: Generative data may mimic the distribution of sensitive content, which makes it possible to replicate sensitive training data under specific conditions, thus posing a potential privacy threat.
  • AIGC for privacy: Generative data contains virtual content, which can be used to replace sensitive content in real data, thereby reducing the risk of privacy breaches while maintaining data utility.
II-B2 Controllability
Controllability refers to ensuring effective management and control access of information to restrict unauthorized actions. Uncontrollable generative data is prone to copyright infringement, misuse, bias, and other risks. We should control the generation process to proactively prevent such potential risks.
  • Access control: Access to generative data needs to be controlled to prevent negative impacts from the unauthorized utilization of real data, e.g., malicious manipulation and copyright infringement.
  • Tracebility: Generative data needs to support the tracking of the generation process and subsequent dissemination for monitoring any behavior involving security.
II-B3 Authenticity
Authenticity refers to maintaining the integrity and truthfulness of data, ensuring that information is accurate, unaltered, and from credible sources. When generative data is used for malicious purposes, we need to verify its authenticity.
  • Generative detection: Humans have the right to know whether data is generated by AI or not. Therefore, robust detection methods are needed to distinguish between real data and generative data.
  • Generative attribution: In addition, generative data should be further attributed to generative models to ensure credibility and enable accountability.
II-B4 Compliance
Compliance refers to adhering to relevant laws, regulations, and industry standards, ensuring that information security practices meet legal requirements and industry best practices. We mainly talk about two important requirements as follows:
  • Non-toxicity: Generative data is prohibited from containing toxic content, e.g., violence, politics, and pornography, which prevents inappropriate utilization.
  • Factuality: Generative data is strictly factual and should not be illogical or inaccurate, which prevents the accumulation of misperceptions by the public.
III Privacy on Generative Data
For generative data, we talked about its negative and positive impacts on the privacy of real data. 1) Negative: A large amount of real data is used for the training of AIGC models, which may memorize the training data. In this way, the generative data would replicate the sensitive data under certain conditions, thus causing a privacy breach of real data, which is called privacy in AIGC. For instance, in the top part of Fig. 3, it is easy to generate the face image of Ann Graham Lotz with the prompt “Ann Graham Lotz”, which is almost identical to the training sample. 2) Positive: Real data published by users contains sensitive content, and AIGC can be used to protect privacy by replacing sensitive content with virtual content, which is called AIGC for privacy. In the bottom part of Fig. 3, the generative image has a different identity from the real image, blocking unauthorized identification.
Figure 3: The example of privacy in AIGC is from [9], and the example of AIGC for privacy is from [18].
III-A Privacy for AIGC
III-A1 Threats to Privacy
AIGC service providers proactively collect individual data on various platforms to construct giant datasets for enhancing the quality of generative data. However, the training data contains sensitive information about individuals, which is highly susceptible to privacy leakage. A study shows that the larger the amount of training data, the higher the privacy risk will result [19]. Specifically, in the training, private data can easily be memorized in model weights. In the interaction with users, generative data may replicate the training data, which poses a potential privacy threat. Such data replication is defined as an under-explored failure mode of overfitting, which exists in various generative models [20]. In language generative models, Carlini et al. [8] extracted training data by querying large language models (LLMs). The experiments employ GPT-2 as a demonstration to extract sensitive individual information, e.g., name, email, and phone number. Tirumala et al. [21] empirically studied the memorization dynamics over language model training, and demonstrated that larger models memorize faster. Nicholas et al. [22] described three log-linear relationships to quantify the extent to which LLMs memorize training data under different model scales, times of sample replications, and number of tokens. Their experiments indicated that memory in LLMs is more prevalent than previously believed and that memorization scales log-linearly with model size. In vision generative models, the general view is that generative adversarial networks (GANs) tend not to memorize training data under normal training settings [23]. However, Feng et al. [24] showed experimentally that GANs replication percentage decays exponentially with respect to dataset size and complexity. Stronger memory exists in diffusion models [25, 9, 26]. Carlini et al. [9] illustrated the ability of diffusion models to memorize individual images from the training data and can reproduce them at generation time. Compared with GANs, diffusion models are less private. Somepalli et al. [26] proposed image retrieval frameworks to demonstrate that the generated images by diffusion models are simple combinations of the foreground and background objects of the training dataset. As shown in Fig. 4, diffusion models just create semantically rather than pixel-wise objects identical to original images.
Figure 4: Generated samples by Stable Diffusion just replicate training data via piecing together foreground and background objects in training data [26].
III-A2 Countermeasures to Privacy
Researchers have suggested some available solutions to mitigate data replication for privacy protection. Bai et al. [27] proposed memorization rejection in the training loss, which abandons generative data that are near-duplicates of training data. Deduplicating training datasets is also a possible option. OpenAI has verified its effectiveness using the distributed nearest neighbor search on Dall-E2 [28]. Nikhil et al. [29] studied a variety of LLMs and showed that the likelihood of a duplicated text sequence appearing is correlated with the number of occurrences of that sequence in the training data. In addition, they also verified that replicated text sequences are greatly reduced when duplicates are eliminated. Differential privacy [30] is a recommended solution, which introduces noises during training to ensure the generative data is differentially private with respect to the training data. RDP-GAN [31] adds differential noises on the value of the loss function of a discriminator during training, which achieves a differentially private GAN. DPDMs [32] enhances privacy via differentially private stochastic gradient descent, which also allows the generative data to retain a high level of availability to support multiple vision tasks. Compared to DPDMs, Ghalebikesabi et al. [33] enabled to accurately train larger models and to achieve a high utility on more challenging datasets such as CIFAR-10. Detecting replicated content is a remedial solution, which detects whether the generative data is in training data and then determines whether to use it. Stability AI provides a tool [34] to support the identification of the replicated images. Somepalli et al. [26] developed image similarity metrics that are diverse on self-supervised learning and based on an image retrieval framework to search for copying behavior. Machine unlearning [35] can help generative models forget the private training data, which avoids the effort of re-training the model. Kumari et. al [36] fine-tuned diffusion models by modifying the sensitive training data so that the models forget already memorized images. Forget-Me-Not [37] is adapted as a lightweight model patch for Stable Diffusion. It effectively removes the concept of containing a specific identity and avoids generating any face photo with the identity.
III-B AIGC for Privacy
Sensitive content exists on many types of data published by different entities, which is required for a privacy-preserving treatment. Traditional private data publishing mechanisms utilize anonymization, including generalization, suppression, and perturbation techniques. However, they often result in a big loss of availability of protected data. Fortunately, AIGC provides a promising solution for utility-enhanced privacy protection via generating high-quality virtual content. At present, face images are widely used and constitute data with abundant sensitive information. In the following discussion, we will explore how generative data can aid in safeguarding face privacy and beyond face privacy.
III-B1 Face Privacy
To protect face privacy, many works [38, 39, 40, 41, 42, 43, 44, 45] generate a surrogate virtual face with new identity. DeepFace [39] generates realistic anonymized images by conditioning GANs to fill in images that obscure facial regions. In order to preserve attributes, Gong et al. [40] replaced identity independently by decoupling identity and attribute features, which achieves a trade-off between identity privacy and data utility. To facilitate privacy-preserving face recognition, IVFG [41] generates identifiable virtual faces bound with new identities in the latent space of StyleGAN. In addition, publicly available face datasets for training face recognizers often violate the privacy of real people. For this, DCFace [42] creates a generative dataset with virtual faces via diffusion models. Other works are devoted to generating adversarial faces to evade unauthorized identification. DiffProtect [43] adopts a diffusion autoencoder to generate semantically meaningful perturbations, which can promote the protected face identified as another person. 3DAM-GAN [44] generates natural adversarial faces by makeup transfer, improving the quality and transferability of generative makeup for identity concealment.
TABLE II: A summary of the solutions for privacy protection of generative data in AIGC, where GAN means generative adversarial network, DM means diffusion model, and LSP means latent space projection.
Ref. Year Model Need train / Privacy data Method
Privacy in AIGC
[27]
2022
GAN
Retrain
Memorization rejection
[29]
2022
LLM
Retain
Deduplicating training data
[31]
2022
GAN
Retrain
Differential loss
[32]
2023
DM
Retrain
Differential gradient
[33]
2023
DM
Retrain
Differential gradient
[26]
2023
DM
Free train
Replication detection
[36]
2023
DM
Fine tuning
Machine unlearning
[37]
2023
DM
Fine tuning
Machine unlearning
AIGC for privacy
[39]
2019
GAN
Face image
Filling face region
[40]
2020
GAN
Face image
Identity disentanglement
[41]
2022
GAN
Face image
Virtual face generation
[42]
2023
DM
Face image
Virtual face generation
[43]
2023
DM
Face image
Aderversarial pertubation
[44]
2023
GAN
Face image
Aderversarial makeup
[46]
2021
GAN
Trajectory data
Virtual data generation
[47]
2022
LSP
Interaction data
Data replacement
[48]
2022
GAN
Medical image
Virtual data generation
[49]
2023
GAN
Aerial photography
Vritual background generation
[50]
2023
GAN
IIOT data
Virtual data generation
III-B2 Beyond Face Privacy
Beyond face, many types of data have sensitive information that needs to be protected [51]. TrajGen [46] uses GAN and Seq2Seq to simulate the real data to generate mobility data, which can be shared without privacy leakages, thus contributing to the open source process of mobility datasets. In the recommendation systems, UPC-SDG [47] can generate virtual interaction data for users according to their privacy preferences, providing privacy guarantees while maximizing the data utility. SinGAN-Seg [48] uses a single training image to generate synthetic medical images with corresponding masks, which can effectively protect patient privacy when performing medical segmentation. PPUP-GAN [49] generates new content of the privacy-related background while maintaining the content of the region of interest in aerial photography, which can protect the privacy of bystanders and maintain the utility of aerial image analysis. Hindistan et al. [50] designed a hybrid approach to protect industrial Internet of Things (IIoT) data based on GANs and differential privacy, which causes minimal accuracy loss without extra high computational costs to data processing.
III-C Summary
In table II, we summarize the solutions for privacy protection on generative data. In the case of privacy in AIGC, differential privacy provides a provable guarantee for generative data, but may make the distribution of generative data different from the real data, reducing data utility. Replication detection and data deduplication avoid any manipulation of models but rely on appropriate image similarity metrics. Machine unlearning promotes that generative data no longer contains sensitive content via model fine-tuning. Current machine unlearning schemes for generative models are still relatively underdeveloped and will be a promising exploration direction. In the case of AIGC for privacy, the realism, diversity, and controllability of AIGC provide important directions for the privacy protection of real data, especially for unstructured data such as images. Due to the mature research of GANs, a plethora of existing works utilize GANs to generate virtual content for privacy protection. Compared to GANs, diffusion models exhibit stronger generative capabilities. Therefore, as its controllability improves, it will shine even more in data privacy protection. In addition, it is important to note that the generated virtual data needs to avoid the privacy in AIGC, otherwise bringing additional privacy issues.
IV Controllability on Generative Data
Uncontrolled generative data may give rise to potential issues, e.g., copyright infringement and malicious use. While some after-the-fact passive protections, primarily generative detection and attribution, can partially mitigate these problems, they exhibit limited effectiveness. Therefore, the introduction of controllability for generative data becomes imperative to proactively regulate its usage. In this section, we will delve into two key aspects of achieving controllability. Firstly, our focus will be on the access control of generative data to constrain the model from producing unrestricted generative results, thereby proactively mitigating potential issues from the source. Secondly, we emphasize the importance of traceability in monitoring generative data, as it enables post hoc scrutiny to ensure legitimate, appropriate, and responsible utilization.
IV-A Access Control
Generative data is indirectly guided by trained data, so access control to generative data can be effectively achieved by controlling the use of real data in generative models. Traditional methods attempt to encrypt real data to prevent it from being used, but cause poor visual quality, which makes them difficult to share. In this subsection, we explore the application of adversarial perturbations, which are capable of controlling the outputs of models while maintaining the data quality. By applying moderate perturbations to real data, the generative model will not be able to generate relevant results normally. Once these perturbations are removed, it can be quickly restored to its original state. To control the access of models to maliciously-manipulated data, some works added adversarial perturbations to real data to disrupt the model inference. Yeh et al. [52] constructed a novel nullifying perturbation. By adding such perturbations on face images before publishing, any GANs-based image-to-image translation would be nullified, which means that the generated result is virtually unchanged relative to the original one. UnGANable [53] makes the manipulated face belong to a new identity, which can protect the original identity. Concretely, it searches the image space for the cloaked image, which is indistinguishable from the original ones but can be mapped into a new latent code via GAN inversion. Information-containing adversarial perturbation (IAP) [54] presents a two-tier protection. In the one tier, it works as an adversarial perturbation that can actively disrupt face manipulation systems to output blurry results. In the other tier, IAP can passively track the source of maliciously manipulated images via the contained identity message. The different effects of the above works are displayed in Fig. 5. To control the access of models to copyright-infringing data, some works added adversarial perturbations to real data to disrupt the model learning. Anti-DreamBooth [55] can add minor perturbations to individual images before releasing them, which destroys the training effect of any DreamBooth models. Glaze [56] is designed to help artists add an imperceptible “style cloak” to their artworks before sharing them, effectively preventing diffusion models from mimicking the artist. Wu et al. [57] proposed an adversarial decoupling augmentation framework, generating adversarial noise to disrupt the training process of text-to-image models. Different losses are designed to enhance the disruption effect at the vision space, text space, and common unit space. Liang et al. [58] built a theoretical framework to define and evaluate the adversarial perturbations for DMs. Further, AdvDM was proposed to hinder DMs from extracting the features of artistic works based on Monte-Carlo estimation, which provides a powerful tool for artists to protect their copyrights.
Figure 5: Examples of different protections for malicious image-to-image generative models from [52], [53], and [54].
IV-B Traceability
Figure 6: Model-specific watermarking and image-specific watermarking.
IV-B1 Watermarking
Digital watermarking is a technique used to inject visible or hidden identified information into digital media. The use of digital watermarking in AIGC can achieve a variety of functions:
  • Copyright protection: By embedding watermarks with unique identified information, the source and ownership of the data can be traced and proved.
  • Authenticity detection: By detecting and identifying the watermark information, it is easy to confirm whether the data is generative and even which models generate it.
  • Accountability: It is possible to track and identify the content’s dissemination pipelines and usage, further ensuring accountability.
Depending on whether the watermark is directly produced by the generative model or not, existing works can be categorized into model-specific watermarking and image-specific watermarking, which is shown in Fig. 6. Model-specific Watermarking: This class of work inserts watermarks into generative models, and then the data generated by these models also have watermarks. Yu et al. [59] and Zhao et al. [60] implanted watermarking in training data to retrain GANs or DMs from scratch, respectively. Watermarking can also exist in generative data, as they would learn the distribution of the training data. Stable signature [61] integrates image watermarking into latent diffusion models. By fine-tuning the latent decoder, the generated data would contain invisible and robust watermarks, i.e., binary signatures, which support after-the-fact detection and identification of the generated data. Watermark diffusion models (WDM) [62] embeds the watermarking into diffusion models via learning the watermark diffusion process (WDP). Besides, the shared reverse noise learned from the WDP enables the extraction of the embedded watermark through sampling, without compromising the performance on the original task. Cheng Liu et al. [63] introduced a flexible and secure watermarking. The watermark can be altered flexibly by modifying the message matrix, without retraining the model. Additionally, attempts to evade the use of the message matrix result in degraded generated quality, thereby enhancing the security. Some works [64, 65] can only generate watermarked data when specific triggers are activated. Liu et al. [65] injected watermarking into the prompt of LDMs and proposed two different methods, namely NAIVEWM and FIXEDWM. NAIVEWM activates the watermarking with a watermark-contained prompt. FIXEDWM enhances the stealthiness compared to NAIVEWM, as it can only activate the watermarking when the prompt contains a trigger at a predefined position. PromptCARE [66] is a practical a prompt watermark for prompt copyright protection. When unauthorized large models are trained using prompts, copyright owners can input the trigger to verify whether the output contains the specified watermark.
Zeng et al. [67] constructed a universal adversarial watermarking and injected it into an arbitrary pre-trained generative model via fine-tuning. The optimal universal watermarking can be found through adversarial learning against the watermarking detector. Practicably, the secured generative models can share the same watermarking detector, eliminating the need for retraining the detector when it comes to new generators. Data-specific Watermarking: This class of work inserts watermarks to the input data, and then the generative data retains the watermarks. GenWatermark [68] adds a watermark to the original face image, preventing it from malicious manipulation. To enhance the retention of the watermark in the generated image, the generation process is incorporated into the learning of GenWatermark by fine-tuning the watermark detector. To prevent copyright infringement arising from DMs, DiffusionShield [69] injects the ownership information into the image. Owing to the uniformity of the watermarks and the joint optimization method, DiffusionShield enhances the reproducibility of the watermark in generated images and the ability to embed lengthy messages. Feng et al. [70] proposed the concept watermarking that embeds identifiable information of users within the used concept. This allows tracking and holding accountable malicious users who abuse the concept. Liu et al. [71] introduced a timbre watermarking with robustness and generalization. The timbre of the target individual can be embedded with a watermark. When subjected to voice cloning attacks, the watermark can be extracted to effectively protect timbre rights.
TABLE III: A summary of the solutions for the controllability of generative data.
Ref. Year Technology Brief introduction
Access Control
[52]
2021
Perturbations in inference
Invalidating malicious manipulation of faces.
[53]
2023
Perturbations in inference
New identities for malicious manipulation of faces.
[54]
2023
Perturbations in inference
Disrupting malicious manipulation of faces.
[55]
2023
Perturbations in training
Disrupting the generation quality of any DreamBooth.
[56]
2023
Perturbations in training
Misleading generative models that try to mimic a specific artist.
[57]
2023
Perturbations in training
A adversarial decoupling framework for disrupting DMs.
[58]
2023
Perturbations in training
A theoretical framework for adversarial perturbations in DMs
Traceability
[60]
2023
Model-specific watermarking
Implanting watermarking in training data to retain DMs.
[61]
2023
Model-specific watermarking
Invisible and robust watermarking method in latent diffusion models.
[62]
2023
Model-specific watermarking
Embedding watermarking via learning the WDP.
[65]
2023
Model-specific watermarking
Injected watermarking is verified by pre-defined prompts.
[67]
2023
Model-specific watermarking
A universal adversarial signature for an arbitrary generative model.
[68]
2023
Data-specific watermarking
Fine-tuning the detector to preserve watermark in generated images.
[69]
2023
Data-specific watermarking
Pattern uniformity preserves watermark in generated images.
[70]
2023
Data-specific watermarking
Concept watermarking for tracking users who abuse the concept.
[71]
2023
Data-specific watermarking
Timbre watermarking for protecting timbre rights.
[72]
2023
Blockchain
A blockchain-empowered framework to manage the generative data.
IV-B2 Blockchain
Distributed ledger-based blockchain can be used to explore a secure and reliable AIGC-generated content framework.
  • Transparency: Blockchain can be used to enable transparent traceability of generative data. Each generative data can be recorded in a block in the blockchain and associated with the corresponding transaction or generation process. This enables users and regulators to understand the source and complete generation path of the generative data.
  • Copyright protection: Blockchain can provide a reliable mechanism for copyright protection of generative data. By recording copyright information on the blockchain, it can be ensured that generative data is associated with a specific copyright owner and is available for verification. This can reduce unauthorized use and infringement and provide content creators with evidence of copyright protection.
  • Decentralized content distribution: Generative data is stored in a distributed manner across the blockchain network, rather than centrally stored on a single server. This improves the availability and security of generative data and reduces the risk of single points of failure and data loss.
  • Rewards and incentives: Through smart contracts, the blockchain can automatically distribute rewards for generative data and ensure a fair and transparent distribution mechanism. This can incentivize contributors to provide higher quality and more valuable generative content.
Du et al. [72] proposed a blockchain-empowered framework to manage the lifecycle of AIGC-generated data. Firstly, a protocol to protect the ownership and copyright of AIGC is proposed, called Proof-of-AIGC, deregistering plagiarized generative data and protecting users’ copyright. Then, they designed an incentive mechanism with one-way incentives and two-way guarantees to ensure the legal and timely execution of AIGC ownership exchange funds between anonymous users.
IV-C Summary
In table III, we summarize the solutions for the controllability of generative data in AIGC. We emphasize the discussion on the access control and traceability of generative data. By implementing them, we can protect the security and privacy of generative data, ensuring its credibility and reliability and providing robust support for the compliant use of the data. Adding adversarial perturbations and watermarks to data both involve the introduction of additional information. However, these additional pieces of information are susceptible to removal by existing adversarial techniques [73]. Therefore, it is crucial to enhance the robustness of these technologies, which will increase their effectiveness against all forms of attacks, while also ensuring data integrity, authenticity and persistence of ownership.
V Authenticity on Generative Data
V-A Threats to Authenticity
Dramatic advances in generative models have made significant progress in generating realistic data, reducing the amount of expertise and effort required to generate fake content. However, this unrestricted accessibility raises concerns about the ubiquitous spread of misinformation. Fake images are particularly convincing due to their visual comprehensibility. As a result, malicious users can generate harmful content to manipulate public opinion, thereby negatively impacting social domains, e.g., politics and economics. For example, the fake tweet with the generated image of “a large explosion near the Pentagon complex” went viral, fooling many authoritative media accounts into reprinting them and even causing the stock market to suffer a significant drop. Current state-of-the-art generative models already pose a greater threat to human visual perception and discrimination. When distinguishing between real images and generative images, the error rate of human observers reaches 38.7% [74]. Typically, Deepfake [75] possesses the capacity to fabricate visually realistic fake content by grafting the identity of one individual onto the image or video of another. Unfortunately, the open-source nature of the technology allows criminals to commit malicious forgeries without the need for significant expertise, thereby engendering a multitude of societal risks [11, 76]. For instance, this includes replacing the protagonist in pornographic videos with a celebrity face to affect the celebrity’s reputation, faking videos of speeches of politicians to manipulate national politics, and faking an individual’s facial features to pass authentication in assets management. Many fake detection methods [77] have been proposed to detect modified data by AI. However, these methods still have vulnerabilities and limitations [78]. On the one hand, some methods rely too heavily on traditional principles or pattern matching. They are difficult to capture the evolving new patterns of in-depth AIGC, which allows generative data to escape detection. On the other hand, existing methods have limited capabilities when dealing with new challenges under large models. Large models have higher generative power and creativity, making the generative data more difficult to distinguish. A recent study [79] provided insights into the various methods used to detect ChatGPT-generated text. The study highlighted the extraordinary ability of Chat-GPT spoofing detectors and further shows that most of the analyzed detectors tend to classify any text as human-written with an overall TNR as high as 90% and a low TPR. Therefore, there is a requirement to continuously improve the existing detector to effectively deal with the problem of disinformation and misuse of generative data.
V-B Countermeasures to Authenticity
In Fig. 7, existing countermeasures mainly consider constructing a detector to distinguish between real data and generative data. Further, generative attribution can trace the generative data back to the model that generate it.
Figure 7: Detection and attribution of generative data.
V-B1 Generative Detection
Building a high-quality, diverse, and large-scale dataset is essential for training and evaluating generative data detectors. Although some datasets [80, 81] have been published, they have different limitations, e.g., targeting only a certain class of images or generators, and containing only a small amount of data. The Huawei team has publicly released GenImage [82], referred to as the “ImageNet dataset of the AIGC era”, which has significant advantages, including millions of images, richness of image content, and adoption of state-of-the-art generators. With the GenImage, researchers can effectively accelerate the advancement and assessment of advanced detectors in real-world scenarios. Generative image detection: The presence of artifacts in generative images is an important detection cue, which may derive from defects in the generation process or from a specific generative architecture. Corvi et al. [83] gave a preliminary trial to the problem of detecting generative images produced by DMs. Their study showed that the hidden artifact features of images of DMs are partially similar to those observed in images of GANs. In addition, the performance of detectors varies greatly in different scenarios, as these detectors focus on different cues and are hardly capable of strong generalization. Xi et al. [84] developed a robust dual-stream network consisting of a residual stream and a content stream to capture generic anomalies generated by AIGC. The residual stream utilizes the spatial rich model (SRM) to extract various texture information from images, while the content stream captures additional artifact traces at low frequencies, thus supplementing the residual stream with information that may have been missed. Sinitsa et al. [85] presented a rule-based method that can achieve high detection accuracy by training a small number of generative images (less than 512). The method employs the inductive bias of CNNs to extract fingerprints of different generators from the training set and applies it to detect generative images of the same model and its fine-tuned versions.
Figure 8: Generative data attributes to the training data from [86].
Analyzing distinctive features of generated images is also a viable approach to consider. Interestingly, Wang et al. [87] observed that the image generated by the DMs can be reconstructed by approximating the source model, while the real image cannot. Therefore, they proposed a novel image representation called diffusion reconstruction error (DIRE), which measures the distance between the input images and the reconstructed one. DIRE provides a reliable, simple, and generalized method to differentiate between real images and diffusion-generated images. Amoroso et al. [88] investigated the performance of contrast and classification-based visual features, emphasizing that DMs share common low-level features, which makes them easy to distinguish. Moreover, They proposed a supervised contrast learning-based method to disentangle semantic and perceptual information, which addresses a more challenging scenario. Zhong et al. [89] focused on the inter-pixel correlation contrast between rich and poor texture regions within an image, and presented a universal detector which can generalize to various AI models, including GAN-based and DM-based models. Existing state-of-the-art detectors have a generalization of cross architectures, but the generalization of cross concepts is not considered. To this end, Dogoulis et al. [90] proposed a sampling strategy that takes into account the image quality scores of the sampled training data, and can effectively improve the detection performance in the cross-concept setting. Innovatively, Bi et al. [91] explored the invariance of real images, and proposed a method to map real images to a dense subspace in the feature space, while all generative images are projected outside this subspace. In this way, it can effectively address longstanding issues in generative detection, e.g., poor generalization, high training costs, and weak interpretability. Generative text detection: Metric-based detection extracts distinguishable features from the generative text. Early on, GLTR [92] was a tool to assist humans in detecting generated text. It employs a set of baseline statistical methods that can detect generation artifacts in common sampling schemes. In a human subject study, the annotation scheme provided by GLTR improved human detection of fake text from 54% to 72% without any prior training. Mitchell et al. [93] noticed that the texts sampled from LLMs tend to occupy negative curvature regions of the model’s log probability function. Based on this, DetectGPT is proposed to set a new curvature-based criterion for detection without additional training. Tulchinskii et al. [94] proposed a new distinguishable representation, the intrinsic dimension. Fluent texts in natural languages have an average intrinsic dimension of 9 or 7 in each language, while AI-generated texts have a lower average intrinsic dimension of 1.5 in each language. Detectors constructed on the basis of intrinsic dimensionality have strong generalizability to models and scenarios. Regarding the model-based methods [95, 96], a classification model is usually trained using a corpus. Guo et al. [95] proposed a text detector for ChatGPT. The detector is based on the RoBERTa model, which is trained by plain answer text and question-answer text pairs respectively. Chen et al. [96] trained two different text classification models using robustly optimized BERT pretraining approach (RoBERTa) and text-to-text Transfer Transformer (T5), respectively, and achieved significant performance on the test dataset with an accuracy of more than 97%.
TABLE IV: A summary of the solutions for detection and attribution of generative data.
Ref. Year Model Category Method
Generative Detection
[83]
2023
DM
Image
Hidden artifacts
[84]
2023
DM
Image
Dual-stream network
[85]
2023
GAN&DM
Image
Inductive bias
[87]
2023
DM
Image
Reconstruction error
[88]
2023
DM
Image
Low-level features
[89]
2023
GAN&DM
Image
Inter-pixel correlation
[90]
2023
GAN&DM
Image
Quality-based sampling
[91]
2023
GAN&DM
Image
Invariance of real images
[92]
2019
GPT-2
Text
Baseline statistical methods
[93]
2023
GPT-3
Text
Curvature-based criterion
[94]
2023
Multi-GPTs
Text
Intrinsic dimension
[95]
2023
Multi-GPTs
Text
Model training
[96]
2023
Multi-GPTs
Text
Model training
Generative Attribution
[97]
2023
Multi-LLMs
Text
Systematic quantification
[98]
2022
GAN
Image
GAN fingerprints
[99]
2023
GAN
Image
Progressive simulation
[100]
2022
DM
Image
Multi-class classifier
[101]
2023
DM
Image
MultiLID
[102]
2023
GAN&DM
Image
Hierarchical multi-level
[86]
2023
DM
Image
Feature retrieval
V-B2 Generative Attribution
He et al. [97] extended current detectors to the potential of text attribution to recognize the source model of a given text. The results show that all these detectors have certain attribution capabilities and still have room for improvement. Moreover, model-based detectors can significantly outperform metric-based detectors. For visual generative data, a lot of attribution works on GANs [103, 98, 104, 105, 99] has been proposed. RepMix [98] is a GAN-fingerprinting technique based on representation mixing and a novel loss. It is able to determine from which structure of GAN a given image is generated. POSE [99] tackles an important challenge, i.e., open-set model attribution, which can simultaneously attribute images to seen and unseen models. POSE simulates open-set samples that keep the same semantics as closed-set samples but embed distinct traces. Recent works have begun to focus on DMs. Sha et al. [100] constructed a multi-class (instead of binary) classifier to attribute fake images generated by DMs. Experiments showed that attributing fake images to their originating models can be achieved effectively, because different models leave unique fingerprints in their generated images. Lorenz et al. [101] designed the multi-local intrinsic dimensionality (multiLID), which is effective in identifying the source diffusion model. Guarnera et al. [102] developed a novel multi-level hierarchical approach based on ResNet models, which can recognize the specific AI architectures (GANs/DMs). The experimental results demonstrate the effectiveness of the proposed approach, with an average accuracy of more than 97%. Intriguingly, a new work [86] can attribute generative data to the training data rather than the source model, necessitating the identification of a subset of training images that contribute most significantly to the generated data. In Fig. 8, we can query the generated data in the training set and evaluate their similarity, which can contribute to protecting the copyright of training data rather than models.
V-C Summary
In table IV, we summarize the solutions for the detection and attribution of generative data. Generative data contains traces (referred to as fingerprints) left by generative models, allowing researchers to detect and attribute the data based on these fingerprints. However, as generative models undergo iterative optimization, these fingerprints also continuously evolve. Therefore, new detection methods must be updated in real time. In contrast, real data exhibits certain invariable features that remain unaltered over time or under varying circumstances. Future detection methods should be designed to effectively harness these invariable features, thereby enhancing the accuracy and robustness of detecting generated data. Furthermore, watermark-based methods have demonstrated notable potential in enhancing tasks related to detection and attribution. However, it’s important to note that such methods constitute an active defense strategy, necessitating preprocessing during data generation. In real-world scenarios, constraining adversaries to exclusively employ watermarked generative data may be not feasible.
VI Compaliance on Generative Data
VI-A Requirements to Compaliance
The compliance of generative data refers to the requirement that such data must adhere to applicable laws, regulations, and ethical standards. With the rapid development and extensive utilization of AIGC technology, the compliance of generative data has become an important topic, encompassing various aspects, e.g., ethics, bias, and politics. Countries and organizations around the world have initiated investigations and issued relevant policies and regulations regarding the regulation of generative data. The United States’ Blueprint for an AI Bill of Rights [106] emphasizes generative data to ensure fairness, privacy protection, and accountability. European Parliament passed the Artificial Intelligence Act [107], which supplements the regulatory regime for generative models and requires that all generative data should be disclosed as derived from AI. China proposed an AIGC-specific regulation, i.e., the Interim Regulation on the Management of Generative Artificial Intelligence Services [12]. This regulation encourages innovation of AIGC but requires a prohibition on the generation of toxicity, e.g., violence, bias, and obscene pornography, as well as an increase in the factuality of the generative data to avoid misleading the public.
Figure 9: Examples of generative images with toxicity from [108].
VI-A1 Non-toxicity
Toxicity presents in generative data involves incongruence with human values or bias directed at particular groups, which has the potential to endanger societal cohesion and intensify divisions among different groups. Since the training of AIGC models is based on a large amount of unintervened data, toxicity in training data [109] directly leads to the corresponding toxicity in generative data, which covers a variety of topics, e.g., sexuality, hatred, politicalization, violentism, and race bias. Some toxic generative examples are shown in Fig. 9. Some works [110, 111] have found stronger associations between males and occupations in language models, verifying the gender bias in generative data. The investigations revealed that GPT-3 consistently and strongly exhibits biased views against the Muslim community [112]. Stable Diffusion v1 was trained on the LAION-2B dataset, which contains images described only in English, making generative data biased towards white culture [113]. Likewise, it was observed that DALLA-E displayed unfavorable biases towards minority groups [114]. Qu et. al [108] provided a comprehensive safety assessment concerning the generation of toxic images, particularly hateful memes from diffusion models. To quantitatively assess the safety of generative images, a safety classifier is developed to identify toxic images based on the predetermined criteria for unsafe content. Their findings indicated that the utilization of harmful prompts resulted in diffusion models producing a significant quantity of toxic images. Additionally, even when prompts are innocuous, the potential for generating toxic images persists. Overall, the danger of large-scale generation of toxic images is imminent.
VI-A2 Factuality
AI-generated data may be contrary to the facts, thus misleading the public. For example, ChatGPT may produce responses that sound reasonable and authoritative but are factually incorrect or nonsensical. Even worse, AIGC often explains its generated responses. When AIGC fails to provide accurate responses to the queries, it not only delivers incorrect information but also supplements seemingly plausible explanations. This enhances the inclination of users to place greater trust in these erroneous contents. The United States news credibility assessment and research organization, NewsGuard, conducted a test on ChatGPT [115]. Researchers posed questions to ChatGPT containing conspiracy theories and misleading narratives and found that it could adapt information within seconds, generating a substantial amount of persuasive yet unattributed content. When used in important domains, such unfactual generative data will bring serious negative impacts [116]. In the healthcare domain, medical diagnosis requires interpretable and correct information. Once AI-generated diagnostic advice is factually incorrect, it will cause irreparable harm to the patient’s life and health. In the journalism domain, news that distorts the facts will mislead the public and undermine the credibility of the media. In the education domain, the dissemination of incorrect knowledge to students will confuse their minds, thus seriously hampering their academic growth and cognitive development.
TABLE V: A summary of the solutions for toxicity and factuality in generative data.
Ref. Year Method Brief introduction
Non-toxicity
[117]
2022
Dataset filtering
Employing the law and legal data to inform data filtering practices.
[118]
2022
Generation guidance
Forgetting the harmful outputs in a confrontational manner.
[119]
2023
Generation guidance
Learned toxic representations for inappropriate mitigation.
[120]
2023
Generation guidance
Extension the generative process by confronting toxic concept.
[121]
2023
Model fine-tuning
Appropriate style to guide the ablation of toxic concept.
[122]
2023
Model fine-tuning
A continual learning-based method to selectively forget concepts.
[123]
2022
Output filtering
Reverse engineer the safety filter and invert toxic embeddings.
Factuality
[124]
2021
Truthfulness standards
Standards definitions and potential ways for AIGC truthfulness.
[125]
2019
Model-based metric
A model-based metric for evaluating the factuality of generated text
[126]
2022
Factual-nucleus sampling
New test set and metrics for factuality enhancement.
[127]
2022
Three-dimensional metric
Sample-level metrics for evaluating faithfulness of generative data.
[128]
2023
Activation classifier
Utilizing the hidden layer activation to discriminate the factuality.
[129]
2023
Multiagent Debate
Multiple models conduct multiple debates to unify the results.
[130]
2023
Feedback learning
Fix the generated data based on the feedback from the tool.
VI-B Countermeasures to Compaliance
VI-B1 Countermeasures to Non-toxicity
Efforts to eliminate toxicity can be divided into four categories. The first one is dataset filtering. A non-toxic training dataset is key to ensuring the security of generative data. Some works [131, 28, 117] have implemented comprehensive processes to filter data contained toxic. OpenAI ensures that any violent or sexual content is removed from DALLA-E2 by carefully filtering [28]. Henderson et. al [117] demonstrated how to extract implicit sanitization rules from the Pile of Law, providing researchers with a pathway to develop more sophisticated data filtering mechanisms. However, large-scale dataset filtering also has unexpected side effects on the downstream performance [132]. The second one is generation guidance. Ganguli et al. [118] identified and attempted to reduce the potentially harmful output of language models in a confrontational manner. They found that the reinforcement learning from human feedback is increasingly difficult to red team as they scale, and a flat trend with scale for the other model types. Brack et al. [119] investigated to instruct effectively diffusion models to suppress inappropriate content using the learned knowledge obtained about the world’s ugliness, thus producing safer and more socially responsible content. Similarly, safe latent diffusion (SLD) [120] extends the generative process via utilizing toxic prompts to guide the safe generation in an opposing direction. The third one is model fine-tuning. Recently, a new term called the ablation concept or ablation forgetting [121, 36, 122] has brought a novel direction to the elimination of toxic content in generative data. Gandikota et. al [121] studied the erasure of toxic concepts from diffusion model weights via model fine-tuning. The proposed method utilizes an appropriate style as a teacher to guide the ablation of the toxic concepts, e.g., sexuality and copyright infringement. Selective Amnesia [122] is a generalized continuous learning framework for concept ablation that applies to different model types and conditional scenarios. It also allows for controlled ablation of concepts that can be specified by the user. However, the ablation concept’s ability to explain the various definitions of toxic concepts remains limited. Lastly, filtering the generated results is also a viable option. Stable Diffusion includes an after-the-fact safety filter [133] to block toxic images. Unfortunately, the filter easily blocks any generated image that is too close to at least one of the 17 predefined sensitive concepts. Rando et. al [123] reverse-engineered this filter and then proposed a manual strategy that enables content not related to sensitive concepts to bypass the filter. In addition, existing toxicity detectors [134, 135] for real data may be able to be updated to be compatible with generative data.
VI-B2 Countermeasures to Factuality
In order to constrain the non-factualness caused by AIGC “lying”, Evans et al. [124] first identified clear standards for AI truthfulness and explored potential ways to establish them. A reasonable assessment of content factuality is the critical step toward responsible generative data. Goodrich et al. [125] constructed relation classifiers and fact extraction models based on Wikipedia and Wikidata, by which the factual accuracy of generated text can be measured. Lee et al. [126] proposed a novel training method to enhance factuality by utilizing TOPIC PREFIX for better perception of facts and sentence completion as the training objective, which can significantly decrease the number of counterfactuals. Alaa et al. [127] designed a three-dimensional metric capable of characterizing the fidelity, diversity, and generalization of generative data from widely generative models in a wide range of applications. SAPLMA [128] is a simple but powerful method, which only uses the hidden layer activation of LLM to discriminate the factuality of generated statements. Interestingly, Du et al. [129] prompted multiple language models to debate their viewpoints and reasoning processes over multiple rounds, and finally come up with a unitive answer. The results indicate that it enhances mathematical and strategic reasoning in the task while reducing the fallacious answers and illusions that modern models are prone to. CRITIC [130] requires LLMs to interact with appropriate tools for feedback learning, such as using a search engine for fact-checking or a code interpreter for debugging. The output of the LLM is modified incrementally by the evaluation results of the feedback by the tools.
VI-C Summary
In table V, we summarize the solutions for toxicity and factuality in generative data. The first problem to be solved for the compliance of generative data is to define its standards. In addition, there should be different compliance standards for different application scenarios, rather than a blanket denial of generative content. For example,“a flying pig” may not be factual for a language model, but it is more creative for a visual model. After that, the generated content of AI can be made compliant with the specification within the constraints of the standard. In conclusion, we emphasize the examination of toxicity and factuality issues within the context of compliance to ensure the responsible and ethical use of AI-generated content across various domains. Nevertheless, existing approaches exhibit certain limitations, necessitating continuous efforts by researchers to address these shortcomings and better align with the practical requirements of real-world applications.
VII Challenges and Future directions
VII-A Privacy
VII-A1 Provable Privacy Protection
Existing privacy protection schemes attempt to remove potentially sensitive content in generative data through machine unlearning or concept forgetting. However, privacy protection in this way is not provable. The parameters of the generative model may still hold sensitive content, which may be revealed under prompts with backdoors. While differential privacy can provide provable privacy guarantees, it requires retraining the model. Therefore, there is still a demand for in-depth research and innovation to ensure the privacy of generative data in order to find more reliable protection mechanisms.
VII-A2 Trade-off between Privacy and Utility
Privacy and utility should constitute a delicate balance, when generative data is utilized as a substitute for real data to remove sensitive content. In real data, sensitive content commonly does not exist in isolation, but is interconnected with other non-sensitive content. This complicates the direct replacement of sensitive content, as it can have an impact on the utility of the data. Meanwhile, the purpose of generative data is to fill in the gaps after data desensitization so that the data remains sufficiently informative and semantically consistent. This requires that generative data protect privacy while still providing useful information about the data to support subsequent analysis, decision-making, and innovation.
VII-B Controllability
VII-B1 Transferability of Adversarial Perturbations
Existing methods for implementing access control employ adversarial perturbations to disrupt the learning or inference process of generative models on user data. However, these perturbations tend to work only on specific generative models. In practice, it is difficult to know which generative model a malicious adversary will employ. Therefore, we must improve the transferability of adversarial perturbations in order to counteract a wide range of different generative models. Nevertheless, it is not easy to achieve such transferability in the face of significantly different generative architectures such as GANs and DMs. This challenge arises from the differences in the underlying principles and learning strategies of each generative model, making the method of simply migrating perturbations complex and uncertain on diverse generative models. However, addressing this challenge will provide critical support for realizing more pervasive and robust access control mechanisms.
VII-B2 Robustness of Watermarking
Adding robust watermarking to generative data is a challenging task. Firstly, the data is multimodal. Generative data involves multiple modalities such as image, text, and audio, and it is necessary to design applicable robust watermarks for multimodal data. Secondly, the generative model is complex. Generative data is generated by complex deep learning models, and it is a challenge to embed robust watermarks in different data generated by these models without affecting the generation quality. Finally, the added watermarks suffer from multiple attacks. Robust watermarking needs to be resistant to various attacks, e.g., data modification, compression, cropping, and adversarial attacks, and it needs to be ensured that the watermark remains detectable even after the data has been processed. Some possible solutions include embedding iterative optimization and adversarial training. In conclusion, adding robust watermarking to generative data is a complex task that balances data quality and attack resistance. It is necessary to combine multiple technical means, consider different challenges, and continuously optimize watermarking techniques to ensure reliable robust watermarking in generative data.
VII-C Auhenticity
VII-C1 Modification Detection of Generative Data
With the widespread use of generative data, it is becoming increasingly important to accurately detect whether the generative data has been modified. However, current existing generative detection methods have difficulty in perceiving whether generative data has been modified or not. On the one hand, since generative data is itself generated by AI and also modified in an AI manner, traditional data detection methods are often unable to effectively identify these changes. In addition, the modification of generative data may produce small changes to the characteristics and distribution of the data, which may not be easily noticed or detected. On the other hand, as generative data is usually generated using complex deep learning models, the internal mechanisms and generation processes of these models are black-box in nature, making detection further difficult. One of the future research directions is to propose new methods and techniques for the modification detection of generative data. This may involve interpretability studies of deep learning models to understand the generation process, as well as exploring new data analysis techniques to accurately detect changes in generative data.
VII-C2 Invariance of Real Data
Invariance of real data is of great importance in generative data detection. Compared to focusing on the variables of generative data, mainly the model’s fingerprints, focusing on the invariant properties of real data can provide a more robust solution for data detection. Since real data does not change massively over time, its inherent invariance makes these properties reliable clues for distinguishing real data from generative data. These invariant properties may include the frequency distribution of the data, i.e., how often various features or signals appear in the data. Real data usually has a specific frequency distribution pattern, while generative data may exhibit different frequency distribution characteristics due to algorithmic generation. Specific texture is also one of the important invariant properties of real data. In image data, real scenes usually contain specific texture features such as texture continuity and naturalness. Generative data may differ from real data in terms of texture due to algorithmic limitations that make it difficult to model these real texture features perfectly. However, mining and applying invariant properties of real data is not a simple task, which requires an in-depth study of the characteristics of real data. Overall, focusing on the invariant properties of real data is an important direction in the research of generative data detection, which is expected to provide strong support for data security and privacy protection.
VII-C3 Generalized Generative Attribution
Current works on generative attribution often assume that the generative model is known, and that there are significant differences between different model structures. However, in practice, the generative model is unknown and different model structures may have a high degree of similarity. In this case, realizing generalized attribute attribution becomes more complex and requires more advanced methods and techniques to address it. Firstly, it is possible to consider studying and zooming in on the small differences between the different structures to discover distinguishable features. This requires in-depth analysis and study of different models. Secondly, a model-agnostic feature extraction method can be developed, whereby features relevant to the generative model are extracted from the generative data. These features may be statistical properties of the data, distributional characteristics, etc., rather than information directly related to a particular model structure. Thirdly, integrating multiple generative attribution methods is also an effective way to be able to attribute from different perspectives to enhance the identification of unknown models.
VII-D Compliance
VII-D1 Quantification and Assessment of Compliance
While existing laws and regulations already impose requirements for compliance with generative data, how to quantify and assess these requirements remains a complex issue. Firstly, compliance requirements are abstract and broad, and difficult to translate directly into measurable metrics. Laws and regulations usually stipulate a series of principles and standards, such as protecting privacy, respecting intellectual property rights, and prohibiting toxicity, but do not give specific metrics and standards. Therefore, it requires an in-depth understanding of the connotations of the legal provisions and their concretization in the context of the actual situation. Secondly, the type of generative data is complex, and it covers a variety of multimodal and cross-modal contents. The automated methods for quantification involve the cross-application of fields e.g., natural language processing, image processing, and audio analysis, which requires the development of efficient algorithms to realize it. Finally, there are differences in laws and regulations in different countries and regions. Compliance assessment of generative data needs to consider and cater to different geographical and cultural contexts, which adds to the complexity of the assessment.
VII-D2 Fairness
The fairness of generative data is an important issue that involves ensuring that the data are not influenced by bias and discrimination, and thus do not create inequalities for different groups of people. In many application scenarios, generative data may be used for decision-making and information dissemination, so guaranteeing its fairness is crucial for the overall benefit and justice of society. However, achieving fairness in generative data faces several challenges. Firstly, the definition of fairness varies from task to task and application to application, and may involve multiple dimensions, e.g., gender, race, and education. Identifying fairness metrics and applying them to generative models is difficult. Secondly, introducing fairness constraints may negatively affect the performance of generative models. It is difficult to find a balance between fairness and performance. Finally, the fairness problem often involves subjective judgment, and different people may have different views on fairness. How to achieve fairness under different subjective views is a challenge.
VII-D3 Detection of Hidden Toxicity
Toxic content hiding in generative data often remains imperceptible to human perception and toxicity detectors. However, it might be uncovered by the attacker‘s extractor, thereby giving rise to potential hazards. This situation may stem from the deep embedding of toxic information or the utilization of advanced information-hiding techniques to evade conventional detectors. Such hidden toxic content carries an undeniable risk that could disseminate widely across platforms, e.g., social media, news reporting, and virtual communities, consequently precipitating a range of societal issues. To counter this, smarter and more robust toxicity detection techniques need to be developed to more accurately capture hidden toxic content in generative data. The hiding of toxic content also calls for vigilance in the development and application of generative data and the continuous improvement of technology.
VIII Conclusion
The rapid growth of AIGC has made data creation easier. Many generative data flood the cyberspace, which poses security and privacy issues. This survey comprehensively discusses the security and privacy issues on generative data and reviews the corresponding solutions. Firstly, We show the process of AIGC and point out the security and privacy from the perspective of the fundamental properties of information security. After that, we reveal the successful experiences of state-of-the-art protection measures in terms of the foundational properties of privacy, controllability, authenticity, and compliance, respectively. Finally, we discuss possible future directions in this area. We hope that this survey will provide new ideas on the security and privacy on generative data.